# Multi-Agents are Social Groups: Investigating Social Influence of Multiple Agents in Human-Agent Interactions

### v1

@@ 2.3. Conversational Agents in Online Discussions
Conversational agents (or ”chatbots”) powered by LLMs such as GPT-4 are capable of holding human-like conversations, and are increasingly being used in group discussion settings. Chatbots have long been used in online discussion scenarios, such as to debate with participants and prompt critical thinking (Tanprasert et al., 2024) or facilitate group discussion by improving discussion efficiency and members’ participation (Hadfi et al., 2023; Kim et al., 2020, 2021). Building on these conversational abilities, chatbots have also been used for persuasive means, to encourage a user or a group of users to alter their attitudes or behaviors. Such systems have been used in health and wellness (Balloccu et al., 2021; Oyebode et al., 2021), education (Ahtinen and Kaipainen, 2020; Tian et al., 2021) and social welfare (Shi et al., 2020). While these systems have demonstrated successful persuasion of their users, researchers have also discussed how design characteristics of the chatbot - such as the perceived identity of the bot (i.e., chatbot or human) (Shi et al., 2020) - can affect its persuasiveness.
@@
Given the widespread application of chatbots for engaging users in real-world scenarios and the growing research interest in this area, the impact of the number of chatbots on persuasive outcomes — specifically the use of multi-agents in persuasion — remains underexplored. Leveraging the potential social influence effects of multi-agent systems may offer a novel means for designers of persuasive systems to enhance their effectiveness: for instance, health and wellness systems (Balloccu et al., 2021; Oyebode et al., 2021) could be designed with multiple agents rather than a single one, to increase the persuasiveness of the intervention. With the increasing accessibility of multi-agent platforms (Section 2.1), designers across various fields are more likely to encounter and develop such systems, making the exploration of their resulting social effects a critical area of investigation.
@@
Given this, we raise the question of whether chatbots and LLM agents are capable of exerting social influence in the same ways that human users do. In particular, we investigate whether the affiliation and accuracy mechanisms work with virtual agents. HCI research regarding the Computers as Social Actors (CASA) phenomenon (Nass et al., 1994) has shown various ways in which humans perceive and act towards machines as though they were human. We extend this line of work by asking: do users feel inclined to agree with a group of virtual agents due to a desire to fit in with them?
@@
This question aims to capture the direct outcomes of human interaction with multiple agents. Specifically, we hypothesize that when agents agree with users, users’ opinions will become more polarized in multi-agent conditions. Conversely, when agents disagree with users, their opinions will shift closer to the agents’ stance. We address this research question using both quantitative and qualitative methods.
@@
To answer this question, we quantitatively test the hypothesis that as the number of agents increases, users will experience stronger social influence. We also use qualitative analysis of open-ended responses to explore the underlying reasons for this social influence. Here, we aim to identify potential mechanisms that drive perceived social influence in interactions with multiple agents and to draw comparisons with mechanisms found in human social group interactions.
@@
Lastly, we explore whether participants with different demographic backgrounds are more easily influenced by multi-agent systems, by analyzing the relationship between demographic factors and the outcomes in RQ1 and RQ2 using regression analysis.
@@ 3. Methods
To understand how single- and multi-agent systems create social influence on humans, we conducted a mixed-methods study comprising an experiment and multiple survey segments, utilizing both quantitative measures and qualitative open-ended questions. This study received approval from our university’s ethics review committee prior to commencement.
Figure 2. Overview of the experiment. Part (a) illustrates the three experimental conditions—1-agent, 3-agent, and 5-agent. Part (b) displays example screenshots of each condition during the self-introduction phase. Part (c) outlines the study procedure, starting with a pre-survey, followed by two rounds of discussions (each followed by a mid-survey), and concluding with a post-survey.
@@ 3.1. Experiment Setup
To investigate how the number of agents influences human opinions, we randomly assigned participants to interact with one, three, or five agents. We chose these numbers based on prior research on persuasion and multi-agent interface designs: on one hand, these numbers represent different group sizes in human communication (Cacioppo and Petty, 1979; Trost et al., 1992); on the other, they reflect a practical range of agent numbers for real-world multi-agent applications (Beinema et al., 2021; Jiang et al., 2023; Park et al., 2023a).
@@
To enhance the generalizability of our findings, we adopted two measures: (1) We chose two social topics for discussion instead of one. Specifically, drawing from recent HCI research on social discussions, we chose the topics “Should self-driving cars be allowed on roads?” (Govers et al., 2024) and “Do violent video games promote violence in youths?” (Yeo et al., 2024). We selected these two topics because they are highly relevant to the everyday lives of our participants, making it easier for them to have opinions and thoughts on them. (2) We designed our agent(s) to have two different stances — agents agreeing with participants on one topic and disagreeing on another — to examine whether alignment or disagreement with participants’ opinions affects the degree of social influence and opinion shift.
@@
The study procedure is shown in Figure 2. Participants first completed a pre-survey that assessed their initial opinions on the social topics and their prior experience related to these topics. They were then randomly assigned to one of three conditions: 1-agent, 3-agent, or 5-agent. In each condition, a host agent appeared to guide the participants through the study. The host agent first introduced the task, then asked all agents and the participant to give a brief self-introduction. Then, for each topic, participants engaged in two rounds of conversation with their assigned agent(s). In each conversation round, the host agent first introduced the topic, then asked the agent(s) or participants to share their opinions on it. After the participants shared their viewpoints, the agent(s) responded to the participants’ statements of the topics and express their own thoughts. After each round of conversations, participants completed a mid-experiment survey to rate their perceived social influence of the agent(s) and their perceptions of the interactions. Once all conversations were finished, participants completed a post-survey to capture their final opinions on the two topics.
@@
3.2. Agent Setup
The agents’ conversations were implemented through a combination of rule-based scripts and the GPT-4 API111gpt-4-1106-preview; https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4. For rule-based scripts, we designed a series of arguments either supporting or opposing the stance for each topic, as detailed in Appendix A.3. These arguments were then crafted into agent dialogue, such as ”I would definitely support [self-driving cars] because I think self-driving cars are great! You can sit back, relax, and let the car do the work for you” or ”I couldn’t agree with [allowing self-driving cars] because I think there are still a lot of issues with self-driving cars that need to be addressed. For example, technical developments are not yet perfect.” Across the three conditions, the same set of arguments was presented: with three and five agents, the different agents took turns to present different arguments; with one agent, the same agent presented all the arguments in the same order (see Figure 3 for an example of the dialogue). This was to ensure that any difference in opinion shifts across the three conditions was not due to differences in content quality or quantity.
We also integrated GPT-4 to enhance agent conversations in two ways: (1) parsing user input, such as extracting the user’s name from their greeting, and (2) generating interactive responses during discussions. When the host agent prompted users to share their opinions on the topics, the agent(s) provided brief feedback based on the user’s inputs, such as by giving a summary of the user’s thoughts or asking follow-up questions. For example, if a user said, ”I agree that self-driving cars should be allowed on the roads”, the next agent might have responded with, ”That’s great to hear! What aspects of self-driving cars do you find most appealing or beneficial?” These responses were tightly controlled through pre-set prompts (Appendix A.4), ensuring that the agent(s) only gave responses relevant to the discussion.
The agents’ avatars and rhetorical styles were designed to be human-like to enhance user acceptance (Sheehan et al., 2020). To avoid the uncanny valley effect (Song and Shin, 2024), we used cartoon-style avatars instead of realistic photos. Each avatar was assigned unique colors to help participants distinguish between the agents. We also ensured gender balance within the 3-agent and 5-agent conditions to minimize the potential effect of agent gender on participants’ opinion changes (Tanprasert et al., 2024).
Figure 4. Screenshot of user interface. The conversation page features an instruction at the top, guiding the user on how to begin the study. Users can enter text in the input field located at the bottom. Once the conversation begins, it is directed by a host agent who provides ongoing guidance throughout.
@@ 3.3. Participants
Participants were recruited via CloudResearch [^2] https://www.cloudresearch.com/. The selection criteria required them to be English speakers and over 18 years old. A total of 104 participants were initially recruited, with two not completing the survey and eight failing the attention check. Ultimately, 94 participants were included in the analysis: 31 participants (F: 17, M: 14) in the 1-agent group, 33 participants (F: 17, M: 15) in the 3-agent group, and 30 participants (F: 15, M: 15) in the 5-agent group. The average ages for each group were as follows: 1-agent group = 38.48 (SD = 12.41), 3-agent group = 38.75 (SD = 12.52), and 5-agent group = 31.87 (SD = 9.46). Participants’ educational backgrounds were distributed as follows: 9 were high school graduates, 16 had some college but no degree, 4 held an associate’s degree, 45 held a bachelor’s degree, 19 held a master’s degree or higher, and 1 preferred not to specify.
The study was conducted in a self-developed online platform (Figure 4); participants were required to complete it on a laptop or desktop computer. The platform’s frontend interface was built using JavaScript and HTML and included three main sections: (1) a login page, (2) an initial attitude questionnaire, and (3) a conversation page for interaction with agents. The study lasted approximately 45 minutes to complete, and each participant was reimbursed US\$5.50.333Following CloudResearch’s payment policy of a minimum of US\$6 per hour: https://go.cloudresearch.com/en/knowledge/setting-up-the-hit-and-payment.
@@ 3.4. Measurements
In this section, we describe the measurements used in this study based on our research questions. Full details of the measurements can be found in Appendix A.2.
@@ 3.4.1. Quantitative
User Opinion (RQ1). We assessed participants’ attitudes toward each topic using five questions adapted from a previous study on social discussions (Hackenburg and Margetts, 2024), such as ”Self-driving cars should be allowed on public roads” and ”Allowing self-driving cars on public roads is a good idea.” These questions were asked twice— before and after the conversations — to measure changes in participants’ stances. Responses were rated on a scale from 1 to 6 (1 = ”Strongly disagree,” 6 = ”Strongly agree”), without a neutral option. This was adapted from previous studies (Chen and Kenrick, 2002; Tanprasert et al., 2024) about manipulating users’ attitudes and was adapted to nudge participants to take a stance. Responses were rated on a scale from 1 to 6 (1 = ”Strongly disagree,” 6 = ”Strongly agree”), without a neutral option. This was adapted from previous studies about manipulating users’ attitudes (Chen and Kenrick, 2002; Tanprasert et al., 2024) and was used to nudge participants to take a stance.
Social Influence (RQ2) To compare the social influence exerted by the agents across different scenarios, we adapted self-reported survey questions from previous research (Kim et al., 2024). We measured two types of social influence - informational influence and normative influence. Informational influence (a change in behavior or attitude based on the belief that others’ information is accurate or reliable) was measured through questions like ”My decision was influenced by the opinion of the agent(s)” capture this type of influence. Meanwhile, normative influence (changes in behavior or attitude driven by the desire to avoid exclusion) was measured using questions like ”During the discussion I felt that I had to agree with the opinion of the agent(s)”.
Control Variables. We also measured factors that could potentially affect perceived social influence. Drawing from previous literature, we asked participants about their domain expertise regarding the topics, with questions such as ”How often do you drive?” and ”Have you ever been in a self-driving car?” These factors can influence individuals’ receptiveness to external influence (Hackenburg and Margetts, 2024). Additionally, we included questions from the AI acceptance scale (Pataranutaporn et al., 2023), as well as measures of participants’ conformity (Mehrabian and Stefl, 1995) and compliance (Gudjonsson, 1989) tendencies.
@@ 3.4.2. Qualitative
To understand reasons for potential opinion change and social influence, we formulated open-ended questions based on social influence theory in human-human interactions. We chose these questions instead of surveys because we found no existing survey that adequately captures social influence in human-agent interactions. Furthermore, given the formative nature of our study, qualitative analysis of these open-ended responses could potentially reveal new insights into different types of human-agent interactions.
User Opinion (RQ1). Participants were asked to articulate their thoughts on the two topics both before and after their conversations with the agent(s). An example question is, ”What are your thoughts on self-driving cars? Do you support or oppose the statement, ’Self-driving cars should be allowed on public roads’? Please explain your reasoning.” This allowed participants to further elaborate on the stances they expressed in the quantitative question.
Social Influence (RQ2). We included three questions to gather participants’ perceptions of the social influence from agent(s) during the discussions. The first was a broad question: ”What do you think of the agent(s) during the discussion?”, which was designed to capture general reactions. The other two focused on key aspects of social influence — accuracy and affiliation (Cialdini and Goldstein, 2004). Specifically, we asked, ”Do you think the arguments presented by the agent(s) were accurate and convincing? Why or why not?” and ”During the conversation, did you feel any pressure to agree with the agent(s)? If so, why?

### v2

@@ 2.3. Conversational Agent in Online Discussion
Chatbots have long been used in online discussion scenarios, such as to debate with participants and prompt critical thinking (Tanprasert et al., 2024) or facilitate group discussion by improving discussion efficiency and member participation (Hadfi et al., 2023; Kim et al., 2020, 2021). Building on its conversational abilities, chatbots have also been used for persuasive means, to encourage a user or group of users to alter their attitudes or behaviours (Li et al., 2025b). Such systems have been used in health and wellness (Balloccu et al., 2021; Oyebode et al., 2021), education (Ahtinen and Kaipainen, 2020; Tian et al., 2021) and social welfare (Shi et al., 2020). While these systems have demonstrated successful persuasion of their users, researchers (Shi et al., 2020) also discussed influencing factors, such as the perceived identity of the chatbot (i.e., chatbot or human) which would affect the outcome of persuasion.
@@
Given the widespread application of chatbots for engaging users in real-world scenarios and the growing research interest in this area, the impact of the number of chatbots on persuasive outcomes, specifically the use of multi-agents in persuasion, remains underexplored. Building on prior literature in social influence within human-human interactions, numerous studies have shown that groups of people, when acting collectively, can exert stronger influence than a single individual (Wilder, 1977; Rutkowski et al., 1983; Spartz et al., 2017)—creating peer pressure (Wang et al., 2016) and encouraging conformity (Gerard et al., 1968) to group norms or opinions. Translating this insight to human–AI interaction, multi-agent systems present a novel opportunity for persuasive system design. For instance, health and wellness systems (Balloccu et al., 2021; Oyebode et al., 2021) could be designed with multiple agents rather than a single one to increase the persuasiveness of the intervention. With the increasing accessibility of multi-agent platforms (Section 2.1), designers across various fields are more likely to encounter and develop such systems, making the exploration of the resulting social effects a critical area of investigation.
@@
Given this, we raise the question of whether chatbots and LLM agents are capable of exerting social influence in the same ways as human users do. In particular, we investigate whether the affiliation mechanism works with virtual agents. HCI research into the Computers as Social Actors (CASA) phenomenon has shown various ways in which humans perceive and act towards machines as though they were human. We extend this line of work by asking: Do users feel inclined to agree with a group of virtual agents due to a desire to fit in with them?
@@
This question aims to capture the direct outcomes of human interaction with multiple agents. Specifically, we hypothesize that when agents agree with users, users’ opinions will become more polarized in multi-agent conditions (i.e., their own opinions will be strengthened more than with a single agent). Conversely, when agents disagree with users, their opinions will shift closer to the agents’ stance. We addressed this research question using both quantitative and qualitative methods.
@@
The first part examines whether users perceive different levels of social influence when interacting with varying numbers of agents. We use quantitative measures to test our hypothesis that as the number of agents increases, users will experience stronger social influence. The second part explores the underlying reasons for this social influence through qualitative analysis of open-ended responses. Here, we aim to identify the potential mechanisms driving perceived social influence in interactions with multiple agents and to draw comparisons with mechanisms found in human social group interactions.
@@
Lastly, we explored which types of participants are more easily influenced by multi-agents by analyzing the relationship between demographic factors and the outcomes in RQ1 and RQ2 using regression analysis. Our goal was to provide more concrete design implications for responsible and inclusive AI systems—ensuring that such technologies do not disproportionately influence or disadvantage particular user groups, and informing future safeguards in the deployment of persuasive AI.
@@ 3. Methods
To understand how single- and multi-agent systems create social influence on humans, we conducted a mixed-methods study combining an experiment and survey, utilizing both quantitative measures and qualitative open-ended questions. This study received approval from our university’s ethics review committee prior to commencement.
Figure 2. Overview of the experiment. Part (a) illustrates the three experimental conditions: 1-agent, 3-agent, and 5-agent. Part (b) displays example screenshots of each condition during the self-introduction phase. Part (c) outlines the study procedure, starting with a pre-survey, followed by two rounds of discussions (each followed by a mid-survey), and concluding with a post-survey.
@@ 3.1. Experiment Setup
To investigate how the number of agents influences human opinions, we randomly assigned participants to interact with one, three, or five agents. We chose these numbers based on prior research on persuasion and multi-agent interface designs: on one hand, these numbers represent different group sizes in human communication (Cacioppo and Petty, 1979; Trost et al., 1992), and on the other hand, they reflect a practical range for real-world multi-agent applications (Beinema et al., 2021; Jiang et al., 2023; Park et al., 2023a; Song et al., 2025).
@@
To enhance the generalizability of our findings, we adapted two approaches: (1) We chose two social topics instead of one. Specifically, drawing from recent HCI research on social discussions, we chose the topics “Should self-driving cars be allowed on roads?” (Govers et al., 2024) and “Do violent video games promote violence in youths?” (Yeo et al., 2024). We selected these two topics as they are closely related to people’s everyday lives, making it easier for participants to have opinions and thoughts on them. (2) We designed two types of agent attitudes—agents agreeing with participants on one topic and disagreeing on another—because we wanted to examine whether alignment or disagreement with participants’ opinions would influence the degree of social influence.
@@
The study procedure is shown in Figure 2. Participants were first presented with an overview of the study, in which they were informed that they would engage in two discussions about social issues (see Figure 3, left). They then completed a pre-survey assessing their initial opinions on the two social topics and their prior experience related to these topics. Participants were randomly assigned to one of three experimental conditions: 1-agent, 3-agent, or 5-agent. In each condition, they were informed that they would be participating in a discussion with a corresponding number of AI agents (see Figure 3, right). Upon entering the conversation interface, a host agent appeared to guide the interaction. The host first introduced the task—discussing two different social topics with AI agents—and then prompted all agents and the participant to provide brief self-introductions. For each topic, participants engaged in two rounds of conversation with the assigned agent(s). In each conversation round, the host agent would first introduce the topic and then ask either agent(s) or participants to share their opinions. After the participants shared their opinions, the agent(s) would respond to the participants’ statements of the topics and express theirs. After each round of conversations, participants completed a mid-survey to rate the social influence and their perceptions of the interactions. The average duration of the two conversations was 23.83 minutes (SD = 9.12). Once all conversations were finished, participants completed a post-survey to capture their final opinions on the two topics.
Figure 3. Study Introduction. (a) Introduction of the study overview, where participants were informed that they would engage in two discussions on social issues. (b) Introduction of the AI-agent conversation, where participants were explicitly told they would be interacting with different numbers of AI agents.
3.2. Agent Setup
The agents’ conversations were implemented through a combination of rule-based scripts and the GPT-4 API111gpt-4-1106-preview; https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4. For rule-based scripts, we designed a series of arguments either supporting or opposing the stance for each topic, as detailed in Appendix A.4. These arguments were then crafted into agent dialogue, such as ”I would definitely support the topic because I think self-driving cars are great! You can sit back, relax, and let the car do the work for you” or ”I couldn’t agree with this topic because I think there are still a lot of issues with self-driving cars that need to be addressed. For example, technical developments are not yet perfect.” Across the three conditions, the same set of arguments was presented: with three and five agents, the different agents took turns to present different arguments; with one agent, the same agent presented all the arguments in the same order (example dialog is shown in Figure 4). This was to ensure that if the three conditions led to different shifts in opinions, it was not because the content quantity presented in each condition varied.
We also integrated GPT-4 to enhance agent conversations in two ways: (1) parsing user input, such as extracting the user’s name from their greeting, and (2) generating interactive responses during discussions. When the host agent prompted users to share their opinions on the topics, the agent(s) would provide brief feedback based on user inputs, such as give a summary or ask a question based on the stage of the conversation. For example, after the user expressed that ”I agree that self-driving cars should be allowed on the roads”, the next agent would say ”That’s great to hear! What aspects of self-driving cars do you find most appealing or beneficial?” These responses were tightly regulated through controlled prompts (Appendix A.5.1), ensuring that the agent(s) only expressed understanding in concise messages, thereby preventing any issues related to AI hallucinations.
The agents’ avatars and rhetorical styles were designed to appear human-like to enhance user acceptance (Sheehan et al., 2020). To avoid the uncanny valley effect (Song and Shin, 2024), we used cartoon-style avatars instead of realistic photos. Each avatar was assigned unique colors to help participants distinguish between the agents. We also ensured gender balance within the 3-agent and 5-agent conditions to minimize the potential effect of agent gender on participants’ opinion change (Tanprasert et al., 2024).
@@ 3.3. Participants
Participants were recruited via CloudResearch [^2] https://www.cloudresearch.com/. The selection criteria required them to be English speakers and over 18 years old. A total of 104 participants were initially recruited, with two not completing the survey and eight failing the attention check. Ultimately, 93 participants were included in the analysis: 31 participants (F: 17, M: 14) in the 1-agent group, 32 participants (F: 17, M: 15) in the 3-agent group, and 30 participants (F: 15, M: 15) in the 5-agent group. The average ages for each group were as follows: 1-agent group = 38.48 (SD = 12.41), 3-agent group = 38.75 (SD = 12.52), and 5-agent group = 31.87 (SD = 9.46). Participants’ educational backgrounds were distributed as follows: 9 were high school graduates, 16 had some college but no degree, 4 held an associate’s degree, 45 held a bachelor’s degree, 19 held a master’s degree or higher, and 1 preferred not to specify.
Figure 5. Screenshot of user interface. The conversation page features an instruction at the top, guiding the user on how to begin the study. Users can enter text in the input field located at the bottom. Once the conversation begins, it is directed by a host agent, who provides ongoing guidance throughout.
The study was conducted in a self-developed online platform (as shown in Figure 5), and participants were required to complete it on a computer. The platform’s frontend interface was built using JavaScript and HTML and included three main sections: (1) a login page, (2) an initial attitude questionnaire, and (3) a conversation page for interaction with agents. The study lasted approximately 45 minutes to complete, and each participant was reimbursed US\$5.50, in line with CloudResearch’s payment policy of a minimum of US\$6 per hour  [^3] https://go.cloudresearch.com/en/knowledge/setting-up-the-hit-and-payment.
@@ 3.4. Measurements
In this section, we describe the measurements used in this study based on our research questions. Further details about the measurements can be found in Appendix A.3.
@@ 3.4.1. Quantitative
User Opinion (RQ1). We assessed participants’ attitudes toward each topic using five questions adapted from a previous study on social discussions (Hackenburg and Margetts, 2024), such as ”Self-driving cars should be allowed on public roads” and ”Allowing self-driving cars on public roads is a good idea.” These questions were asked twice—before and after the conversations—to measure changes in participants’ stances. Responses were rated on a scale from 1 to 6 (1 = ”Strongly disagree,” 6 = ”Strongly agree”), without a neutral option. This follows previous studies (Chen and Kenrick, 2002; Tanprasert et al., 2024) about manipulating users’ attitudes and was adapted to nudge participants to take a stance.
Social Influence (RQ2). To compare the social influence exerted by the agents across different scenarios, we adapted self-reported survey questions from previous research (Kim et al., 2024). There are two types of social influence measured - informational influence and normative influence. Informational influence refers to a change in behavior or attitude based on the belief that others’ information is accurate or reliable. For example, questions such as ”My decision was influenced by the opinion of the agent(s)” capture this type of influence. In contrast, normative influence reflects changes in behavior or attitude driven by the desire to avoid exclusion. An example item for this influence is ”During the discussion I felt that I had to agree with the opinion of the agent(s)”.
Control Variables. We also inquired about factors that could potentially affect the results of social influence. Drawing from previous literature, we asked participants about their domain expertise regarding the topics, with questions such as ”How often do you drive?” and ”Have you ever been in a self-driving car?” These factors can influence individuals’ receptiveness to external influence (Hackenburg and Margetts, 2024). Additionally, we included questions from the AI acceptance scale (Pataranutaporn et al., 2023), as well as measures of conformity (Mehrabian and Stefl, 1995) and compliance (Gudjonsson, 1989) tendencies. While these factors were not influenced by the agent(s) settings, they represent inherent characteristics of the participants that could potentially moderate the results.
@@ 3.4.2. Qualitative
To understand reasons for potential opinion change and social influence, we formulated open-ended questions based on social influence theory in human-human interactions. We chose open-ended questions instead of surveys because no existing survey adequately captures social influence in human-agent interactions, and open-ended questions can reveal potential differences between agent(s) and human interactions.
User Opinion (RQ1). Participants were asked to articulate their thoughts on the two topics both before and after their conversations with the agent(s). An example question is, ”What are your thoughts on self-driving cars? Do you support or oppose the statement, ’Self-driving cars should be allowed on public roads’? Please explain your reasoning.” This was designed to complement the findings from the quantitative results in RQ1.
Social Influence (RQ2). We included three questions to gather participants’ perceptions of the social influence from agent(s) during the discussions. The first was a broad question: ”What do you think of the agent(s) during the discussion?”, which was designed to capture general reactions. The other two focused on key aspects of social influence—accuracy and affiliation (Cialdini and Goldstein, 2004). Specifically, we asked, i.e., ”Do you think the arguments presented by the agent(s) were accurate and convincing? Why or why not?” and ”During the conversation, did you feel any pressure to agree with the agent(s)? If so, why?